Round to nearest cent or hundredth percent as needed: (Do not round intermediate calculations. Round your answer to the nearest cent.) Principal ?   Rate 11%   Time 2 months Simple Interest $1,250
Answer #1

Given ,

Simple interest ( I) = 1,250

Rate ( R )= 11% p.a i.e 11/100

Time ( T ) = 2 months i.e 2/12 = 1/6

P= I / ( R × T )

= 1250 × 100 × 6 / 11× 1

= 750,000 / 11

P = $68,182 ( answer )

Note :-

Months = 2

In years = 2/ 12

= 1/6

So , The Principal amount is $ 68,182

If you found this answer useful please give it a uplike .

Thankyou !!!